which of the following is the solution to the equation 2(x-10)=3x-4?
a.) 1
b.) -1
c.) -6
d.) -16

Answer by Edwin McCravy(20056)

2(x - 10) = 3x - 4

Let's find out if the answer is 

a.) 1

by substituting 1 for x in

2(x - 10) = 3x - 4
2(1 - 10) = 3(1) - 4
    2(-9) = 3 - 4
      -18 = -1

No that's false. So the answer is not a.) 1
----------------------------------------------

Now let's find out if the answer is 

b.) -1

by substituting -1 for x in

2(x - 10) = 3x - 4
2(-1 - 10) = 3(-1) - 4
    2(-11) = -3 - 4
      -22 = -7

No that's false. So the answer is not b.) -1
----------------------------------------------

Now let's find out if the answer is 

c.) -6

by substituting -6 for x in

2(x - 10) = 3x - 4
2(-6 - 10) = 3(-6) - 4
    2(-16) = -18 - 4
      -32 = -22

No that's false. So the answer is not c.) -6
----------------------------------------------

Now let's find out if the answer is 

d.) -16

by substituting -16 for x in

2(x - 10) = 3x - 4
2(-16 - 10) = 3(-16) - 4
    2(-26) = -48 - 4
      -52 = -52

Yes that's true! So the answer is d.) -16

Edwin